# CoMaTrack: Competitive Multi-Agent Game-Theoretic Tracking with Vision-Language-Action Models

### 24 Mar 2026 v1

@@ Abstract
Embodied Visual Tracking (EVT), a core dynamic task in embodied intelligence, requires an agent to precisely follow a language-specified target. Yet most existing methods rely on single-agent imitation learning, suffering from costly expert data and limited generalization due to static training environments. Inspired by competition-driven capability evolution, we propose CoMaTrack, a competitive game-theoretic multi-agent reinforcement learning framework that trains agents in a dynamic adversarial setting with competitive subtasks, yielding stronger adaptive planning and interference-resilient strategies. We further introduce CoMaTrack-Bench, the first benchmark for competitive EVT, featuring game scenarios between a tracker and adaptive opponents across diverse environments and instructions, enabling standardized robustness evaluation under active adversarial interactions. Experiments show that CoMaTrack achieves state-of-the-art results on both standard benchmarks and CoMaTrack-Bench. Notably, a 3B VLM trained with our framework surpasses previous single-agent imitation learning methods based on 7B models on the challenging EVT-Bench, achieving 92.1% in STT, 74.2% in DT, and 57.5% in AT. The benchmark code will be available at https://github.com/wlqcode/CoMaTrack-Bench
@@ 1 Introduction
  • First fusion of multi-agent competitive game theory and RL in EVT: We devise a co-evolving adversarial training loop where opponent strategies dynamically escalate task difficulty. Crucially, a compact 3B-parameter VLM trained under CoMaTrack surpasses all existing single-agent methods leveraging 7B-parameter models, confirming that strategic competition—not model scale—drives robust generalization.
  • Open-sourcing CoMaTrack-Bench: We release the first multi-agent adversarial benchmark for EVT, featuring dynamic dueling scenarios with adaptive opponents to enable rigorous, real-world-aligned evaluation of tracking robustness—addressing the idealized limitations of prior benchmarks.
@@ 2.2 Embodied Visual Tracking
This imitation-centric foundation, though effective in constrained settings, faces inherent ceilings. Expert data collection is costly, biased toward ideal trajectories, and IL policies lack exploration mechanisms. In contrast, this work presents the first multi-agent competitive RL paradigm for EVT task.
@@ 2.3 Reinforcement Learning in VLN
We present the first RL-based framework for EVT, and further move beyond single-agent RL by formulating tracking as a multi-agent game-theoretic training process. By leveraging competitive interactions, agents become each other’s adaptive environment, creating an automatically escalating curriculum that is difficult to obtain with static reward designs or offline data alone.
@@ 3.6 CoMaTrack Benchmark
To systematically evaluate tracking robustness under active adversarial interactions and bridge the realism gap in existing benchmarks, we introduce CoMaTrack-Bench, the first competitive multi-agent benchmark for EVT. Unlike prior benchmarks that evaluate tracking under idealized or passively disturbed conditions, CoMaTrack-Bench features dynamic dueling scenarios where a tracker agent must pursue a language-specified target while contending with adaptive opponents that actively interfere with the tracking task.
@@ 4.1 Data Collection
Question Answering Data Organization. To equip the model with real-world 3D spatial reasoning, we incorporate ScanQA [scanqa]. We further leverage LLaVA-Pretrain [llava] to enhance general visual-language understanding. To broaden and diversify representation learning across wider domains, we additionally incorporate SYNTH-PEDES [SYNTH-PEDES]. Finally, we employ RefCOCO [refcoco] and Flickr30k [flickr30k] to jointly strengthen grounded language understanding and image captioning capabilities.
Multi-Task Navigation Data Collection. We collect multi-task navigation data from diverse sources. First, we incorporate instruction-following data from YouTube videos like NaViLA [navila] and use a shortest-path follower in the Habitat simulator [habitat3] to generate expert action sequences for R2R-CE [r2r] and RxR-CE [rxr]. Second, we collect object-goal navigation data from HM3D ObjectNav [hm3d] and HM3D OVON [ovon].

### 31 Mar 2026 v2

@@ Abstract
Embodied Visual Tracking (EVT), a core dynamic task in embodied intelligence, requires an agent to precisely follow a language-specified target. Yet most existing methods rely on single-agent imitation learning, suffering from costly expert data and limited generalization due to static training environments. Inspired by competition-driven capability evolution, we propose CoMaTrack, a competitive game-theoretic multi-agent reinforcement learning framework that trains agents in a dynamic adversarial setting with competitive subtasks, yielding stronger adaptive planning and interference-resilient strategies. We further introduce CoMaTrack-Bench, the first open-source Habitat-based benchmark protocol and episode set for language-conditioned competitive EVT featuring dynamic dueling, featuring game scenarios between a tracker and adaptive opponents across diverse environments and instructions, enabling standardized robustness evaluation under active adversarial interactions. Experiments show that CoMaTrack achieves state-of-the-art results on both standard benchmarks and CoMaTrack-Bench. Notably, a 3B VLM trained with our framework surpasses previous single-agent imitation learning methods based on 7B models on the challenging EVT-Bench, achieving 92.1% in STT, 74.2% in DT, and 57.5% in AT. The benchmark code will be available at https://github.com/wlqcode/CoMaTrack-Bench
@@ 1 Introduction
  • First integration of competitive multi-agent game-theoretic RL with modern VLA architectures for language-conditioned embodied visual tracking: We devise a co-evolving adversarial training loop where opponent strategies dynamically escalate task difficulty. Crucially, a compact 3B-parameter VLM trained under CoMaTrack surpasses all existing single-agent methods leveraging 7B-parameter models, confirming that strategic competition—not model scale—drives robust generalization.
  • Open-sourcing CoMaTrack-Bench: We release the first open-source benchmark suite that standardizes competitive EVT evaluation in Habitat with language instructions, featuring dynamic dueling scenarios with adaptive opponents to enable rigorous, real-world-aligned evaluation of tracking robustness—addressing the idealized limitations of prior benchmarks.
@@ 2.2 Embodied Visual Tracking
Beyond imitation-centric pipelines, reinforcement learning and game-theoretic approaches have long been explored for active visual tracking. Pioneering work such as AD-VAT [advat] introduced an asymmetric dueling mechanism where tracker and target agents engage in adversarial competition, demonstrating that co-evolutionary training yields more robust policies than single-agent paradigms. This was further theoretically analyzed in AD-VAT+ [adcat+] and extended to distraction robust tracking via mixed cooperative-competitive multi-agent games [ts]. Subsequent studies validated adversarial training for tracking generalization across diverse environments [zhong2023rspt, zhong2024empowering]. However, these earlier frameworks primarily operated in non-language-conditioned settings with traditional perception backbones. HIEVT [wu2025hievt] proposes a hierarchical instruction-aware framework using LLM-based goal alignment, while  [wu2025vlm] introduces a self-improving VLM framework for embodied tracking with memory-augmented reflection mechanisms. In contrast, CoMaTrack adapts the competitive game-theoretic RL paradigm to modern VLA models, explicitly incorporating natural language instructions and leveraging large-scale VLM priors to enable robust, instruction-aware tracking under strategic adversarial interference.
@@ 2.3 Reinforcement Learning in VLN
We present the first GRPO-based online RL fine-tuning framework for VLA models in language-conditioned EVT, and further move beyond single-agent RL by formulating tracking as a competitive multi-agent game-theoretic training process. By leveraging competitive interactions, agents become each other’s adaptive environment, creating an automatically escalating curriculum that is difficult to obtain with static reward designs or offline data alone.
@@ 3.6 CoMaTrack Benchmark
To systematically evaluate tracking robustness under active adversarial interactions and bridge the realism gap in existing benchmarks, we introduce CoMaTrack-Bench, the first open-source benchmark suite that standardizes competitive EVT evaluation in Habitat with language instructions. Unlike prior benchmarks that evaluate tracking under idealized or passively disturbed conditions, CoMaTrack-Bench features dynamic dueling scenarios where a tracker agent must pursue a language-specified target while contending with adaptive opponents that actively interfere with the tracking task.
@@ 4.1 Data Collection
Question Answering Data Organization. To equip the model with real-world 3D spatial reasoning, we incorporate ScanQA [scanqa]. We further leverage LLaVA-Pretrain [llava] to enhance general visual-language understanding. To broaden and diversify representation learning across wider domains, we additionally incorporate SYNTH-PEDES [SYNTH-PEDES]. Finally, we employ RefCOCO [refcoco] and Flickr30k [flickr30k] to jointly strengthen grounded language understanding and image captioning capabilities. Multi-Task Navigation Data Collection. We collect multi-task navigation data from diverse sources. First, we incorporate instruction-following data from YouTube videos like NaViLA [navila] and use a shortest-path follower in the Habitat simulator [habitat3] to generate expert action sequences for R2R-CE [r2r] and RxR-CE [rxr]. Second, we collect object-goal navigation data from HM3D ObjectNav [hm3d] and HM3D OVON [ovon].